OpenStudy (anonymous):

Assume the reserve requirement is 10%. First National Bank receives a deposit of $5,400. If there is no slippage, how much could the money supply expand? Explain and show your work. I just want to know how would i set up the problem in this question.

jimthompson5910 (jim_thompson5910):

Not 100% sure, but I think you just set up the ratio 10/100 = 5400/x then solve for x 10/100 = 5400/x 10x = 5400*100 10x = 540000 x = 540000/10 x = 54000 So they could release $54,000 into the money supply if they received a deposit of $5,400

OpenStudy (anonymous):

Thank you, yeah I am confuse to and am not sure my teacher want the math or just explain, that why i ask on here how to do it, it for my econ class. thank you again

jimthompson5910 (jim_thompson5910):

yw, I would double check with another econ student since econ isn't my best strength

OpenStudy (anonymous):

thanks is there econ on here?

jimthompson5910 (jim_thompson5910):

yeah i think so, look in the "find more subjects tab" or you might find someone in econ in the math section, but idk

OpenStudy (dwelsher):

4860?